OpenStudy (anonymous):

Which expression is equivalent to six raised to the fourth power? four raised to the sixth power two raised to the fourth power times three three raised to the eighth power thirty-six times six raised to the second power

OpenStudy (anonymous):

6^4 4^6 2^4 * 3 3^8 36*6^2

OpenStudy (anonymous):

which one is it?

OpenStudy (anonymous):

I was typing them out so I can see it visually. 6^4=1296

OpenStudy (anonymous):

It's 36*6^2. 6^2=36. 36*26=1296

OpenStudy (anonymous):

You could use a calculator

OpenStudy (anonymous):

Ohh ok haha, Thanks!